I need help with this: x^2=18x - 81 Thanks
Answer by ankor@dixie-net.com(22740)

x^2 = 18x - 81
arrange as a quadratic equation
x^2 - 18x + 81 = 0
A perfect square which factors to
(x-9)(x-9) = 0
x = 9
:
;
Check in original equation, x=9
9^2 = 18(9) - 81
81 = 162 - 81